# Linguistics and Human Brain: A Perspective of Computational Neuroscience

### v1

@@ 3 Computational Neuroscience Methods for Linguistics
Computational neuroscience provides a methodological bridge between formal linguistic theory and empirical neural data. As illustrated in Fig. 2, it transforms abstract linguistic structures into computational representations that can be quantitatively compared with neural activity through a set of interrelated tools and analytical frameworks. This section introduces the core components of this bridge. We first outline the methodology of computational neuroscience, and then describe four technical pillars: Measurement Methods (neural activity recording), Word Embeddings (numerical representations of language), Evolving Language Models, and Neural Coding (analytical frameworks that link model representations to brain signals). Finally, we present an integrative case study to illustrate how these components operate jointly and support model–brain alignment research in linguistics and neuroscience.
@@ 3.2 Brain Activity Measurement for Linguistic Tasks
Testing linguistic theories within a computational framework requires reliable measurements of neural activity during language processing. Research on the neural basis of language therefore employs multiple techniques that capture neural responses across spatial, temporal, and representational dimensions [99, 75, 236, 113, 84]. Because language processing emerges from dynamic interactions across distributed brain regions and multiple scales, no single method is sufficient. Instead, measurement techniques are selected and combined according to the research focus, such as spatial localization, temporal resolution, or representational specificity.
@@
    The main advantage of fMRI is its high spatial resolution, enabling accurate localization of cortical and subcortical regions involved in language processing, including Broca’s area, Wernicke’s area, and the angular gyrus. It also distinguishes spatial activation patterns associated with different linguistic levels such as phonetic, lexical, and syntactic processing [18, 195, 19, 241, 256]. Its primary limitation is low temporal resolution: the BOLD response is delayed by approximately 4–8 seconds, with signal peaks often occurring 6–8 seconds after neuronal activation. Consequently, millisecond-scale processes such as rapid semantic integration cannot be directly captured [80, 154]. In addition, measurements are sensitive to physiological noise such as cardiac and respiratory activity, which complicates experiments involving infants or patients with language disorders. Overall, fMRI provides high spatial but limited temporal resolution [16].
@@
    EEG (Electroencephalography): EEG records weak electric field fluctuations generated by synchronous firing of large neuronal populations using scalp electrode arrays [140]. Its key advantage is millisecond-level temporal resolution, allowing precise tracking of rapid neural dynamics during language processing, including responses related to semantic and syntactic analysis. However, electrical signals are attenuated and spatially blurred when passing through brain tissue, skull, and scalp, resulting in poor spatial localization and difficulty in identifying precise neural sources [172, 152]. EEG signals are therefore highly sensitive to synchronized cortical activity but provide limited spatial precision [149].
@@
    MEG (Magnetoencephalography): MEG measures weak magnetic fields generated by neuronal electrical activity using Superconducting Quantum Interference Devices (SQUIDs) [44, 90]. Synchronous neuronal currents produce magnetic fields that can be detected outside the head, providing indirect measurements of neural activity. Like EEG, MEG offers millisecond temporal resolution and can track rapid processing stages from phonetic perception to syntactic and semantic integration.
    Compared with EEG, MEG generally provides better spatial localization because magnetic fields are minimally distorted by skull and scalp tissues. Source localization based on magnetic field distributions therefore reduces uncertainties associated with EEG inverse solutions [92]. This combination of high temporal resolution and improved spatial accuracy makes MEG particularly suitable for studies requiring precise spatiotemporal characterization of language processes, such as phonetic-to-lexical conversion and syntactic processing [208, 210].
@@
    ERP (Event-Related Potentials): ERP is not an independent measurement technique but an analysis method derived from EEG recordings [149]. EEG signals are time-locked and averaged across repeated presentations of language stimuli, allowing stimulus-related neural responses to be extracted while reducing spontaneous activity and noise [128]. ERP therefore retains EEG’s millisecond temporal resolution and provides fine-grained information about temporal stages of language processing.
    Many ERP components have been identified in language research, among which N400 and P600 are the most widely studied. The N400 component, typically peaking around 400 ms after stimulus onset, is strongly associated with semantic processing; its amplitude increases when stimuli are semantically incongruent or unexpected. The P600 component, peaking around 600 ms, is primarily linked to syntactic processing and increases when sentences contain syntactic violations or elevated structural complexity [229, 178, 127].
@@
    fNIRS (functional Near-Infrared Spectroscopy): fNIRS is a non-invasive optical imaging technique that measures brain activity through changes in tissue absorption of near-infrared light associated with blood oxygenation [197, 235]. Near-infrared light emitted from scalp-mounted optodes penetrates superficial tissues and is differentially absorbed by oxyhemoglobin and deoxyhemoglobin. Detectors capture diffusely reflected light, and concentration changes are estimated using the Beer–Lambert law, allowing indirect inference of neural activity. Like fMRI, fNIRS relies on hemodynamic responses associated with neural activation.
    fNIRS offers several practical advantages. The equipment is portable and does not require shielded environments, enabling experiments in more natural communication settings. It is also less sensitive to motion artifacts than fMRI and EEG, making it suitable for studies involving infants, children, or patients with language disorders, as well as tasks involving speech or facial movements [242]. Its temporal resolution lies between fMRI and EEG/MEG, allowing tracking of sub-second hemodynamic changes during tasks such as lexical recognition and sentence comprehension. Spatial resolution typically reaches the centimeter scale, sufficient for localizing cortical language areas [71, 211].
    Limitations include lower spatial resolution than fMRI and restricted penetration depth, which prevents reliable measurement of deep subcortical structures. Optical signals are also influenced by individual physiological factors such as scalp thickness, hair properties, and superficial blood flow, requiring careful experimental design to ensure consistency across participants [137, 191]. In addition, the limited number of channels restricts whole-brain coverage, making fNIRS more suitable for targeted cortical studies [211].
@@
    ECoG (Electrocorticography): ECoG is an invasive neurophysiological technique that records electrical potentials directly from the exposed cortical surface using subdural electrode grids or strips placed intraoperatively or via craniotomy [100]. Unlike scalp EEG, ECoG avoids signal attenuation and distortion from the skull and scalp, yielding high signal amplitude, excellent spatial resolution, and sub-millisecond temporal precision [112]. These characteristics make it uniquely capable of resolving fine-grained, layer-specific and region-specific neural dynamics underlying language processing, including articulatory planning, phonemic encoding, lexical access, and syntactic binding. ECoG is widely used in presurgical mapping for epilepsy or brain tumor patients, allowing direct localization of essential language areas such as Broca’s region, superior temporal gyrus, and inferior parietal cortex with high clinical and scientific reliability [53]. High-frequency broadband power modulations in ECoG are particularly sensitive to local neuronal population activity and correlate strongly with task-specific language computations. Key limitations include its invasive nature, restricted clinical eligibility, limited spatial coverage determined by electrode placement, and inherent ethical constraints in healthy participants.
@@
    SEEG (Stereo‐Electroencephalography): SEEG is an invasive three‐dimensional stereotactic technique that uses multiple thin, depth electrodes implanted orthogonally through small burr holes to record electrical activity from deep brain structures, white matter tracts, and circumscribed cortical regions [263]. Unlike ECoG, which samples the cortical surface, SEEG targets subcortical nuclei, insular cortex, cingulate gyrus, medial temporal regions, and other deep areas critical for language control, semantic memory, and pragmatic integration [6]. SEEG combines sub-millisecond temporal resolution with precise three‐dimensional localization, enabling dissection of spatiotemporal networks involved in both overt and covert language processing [269]. It is extensively used in epilepsy presurgical evaluation to map language-related networks and identify epileptogenic zones without compromising essential functions. Similar to ECoG, SEEG is restricted to clinical populations, involves surgical risk, and has limited sampling density across the whole brain [183]. Nevertheless, its ability to record from deep and hidden structures makes SEEG irreplaceable for investigating subcortical‐cortical circuits supporting language comprehension, production, and monitoring.
3.3 Word Embedding
@@
Numerical representations of language form a central link between linguistics, computer science, and neuroscience. Word embedding, also known as lexical vector representation, has become a foundational technique in Natural Language Processing (NLP) and in studies of neural language mechanisms [158]. The central idea is to map discrete words into a continuous high-dimensional vector space, where semantic similarity is reflected in geometric relations among vectors. This mapping enables symbolic language to be processed computationally. Unlike traditional one-hot encoding, which only indicates word identity and ignores semantic relations, word embeddings encode words through distributed numerical features that capture semantic associations and significantly improve language modeling and analysis [93].
@@
Word embeddings are typically learned using machine learning or deep learning models and are grounded in the distributional semantic hypothesis, which states that word meaning is shaped by contextual usage patterns [207]. Based on training strategies and model architectures, mainstream approaches fall into two broad categories. The first includes static embedding models such as Word2Vec and GloVe. Word2Vec learns word vectors by predicting local context through Skip-gram or CBOW architectures, while GloVe combines global word frequency statistics with local co-occurrence information. Both methods generate fixed vectors that capture semantic similarity and analogical relations between words [186, 40]. The second category consists of context-dependent embeddings produced by modern Transformer-based pre-trained language models. These models dynamically generate representations conditioned on context, addressing polysemy that static embeddings cannot resolve and improving semantic representation quality [120, 1]. All such approaches rely on large-scale unlabeled corpora and learn word co-occurrence patterns that place semantically related words close in vector space, providing effective foundations for downstream language tasks.
@@
Word embedding plays a central role for following three reasons. First, it connects computational linguistics with neurolinguistics [132]. The distributional principle underlying embeddings parallels the brain’s distributed representation of lexical semantics, and similarities in embedding space often correspond to similarities in neural activity patterns observed with techniques such as fMRI and MEG [162]. Understanding embedding representations therefore facilitates comparison between computational and neural semantic structures. Second, embeddings form the numerical input foundation of modern language models. Natural language, being symbolic and discrete, must be transformed into continuous numerical representations before processing by neural networks. From early static embeddings to context-dependent representations in modern LLMs, embedding quality strongly influences model performance [58]. Since later sections analyze deep learning models for neural simulation, reviewing embedding principles clarifies how linguistic information enters computational models. Third, research on extending embeddings to complex expressions informs studies of neural representations of higher-level linguistic structure [8]. This work emphasizes not only lexical processing but also sentence and discourse-level representations. The strengths and weaknesses of existing extension methods, particularly regarding syntactic preservation and semantic robustness, mirror challenges in modeling hierarchical neural language processing. Reviewing these methods therefore provides conceptual guidance for analyzing how the brain encodes complex linguistic structures.
In summary, word embedding and its extensions provide the core framework for numerical language representation. They have driven rapid progress in NLP while also offering essential tools and theoretical insights for interdisciplinary studies of neural language mechanisms.
@@ 3.4 The Evolving Language Models
Language models are central tools at the intersection of computational neuroscience and linguistics. They learn statistical regularities from textual data, extract semantic information, and generate outputs consistent with human language conventions. Their fundamental function is to predict lexical sequences based on preceding or surrounding context.
Research on language models is important both theoretically and practically. Theoretically, their development parallels attempts to understand neurocognitive mechanisms underlying language processing. Evaluating how models capture contextual dependencies and hierarchical structure provides quantitative references for studying operations of the brain’s language network. Practically, language models form the backbone of many NLP applications, including machine translation, text generation, semantic analysis, etc. Improving their alignment with human cognition is therefore considered an important step toward more general artificial intelligence systems.
@@
    PCFG: Probabilistic Context-Free Grammars explicitly model hierarchical syntactic structure and therefore can, in principle, capture long-distance dependencies. However, both PCFG and n-gram approaches suffer from limited generalization, as they cannot exploit semantic similarity across words or structural similarity across contexts. As a result, they have largely been replaced by neural network approaches that learn distributed representations and achieve better generalization performance [110].
@@
    RNN: Recurrent Neural Networks represent the first generation of neural language models. They encode contextual information by mapping words into embedding spaces and updating hidden states sequentially as new inputs are processed [67]. Although RNNs do not explicitly encode hierarchical structure, their recurrent updates allow them to capture contextual dependencies, including long-distance relations. Long Short-Term Memory (LSTM) networks are a widely used variant that mitigate gradient vanishing through gating mechanisms, substantially improving long-range dependency modeling and dominating early neural language modeling systems [102].
@@
    • Autoregressive LLMs: These models generate text sequentially from left to right by predicting the next token given preceding context. The GPT series is the most prominent example, with parameter scales growing from millions to trillions and training corpora spanning diverse web and literary sources [199, 1, 26, 200, 179]. Such models excel at text generation, dialogue, and long-form continuation tasks. Related systems include ERNIE Bot and the GLM family. Their primary strength lies in fluent generation, although strictly unidirectional modeling limits full bidirectional context use [228, 63, 267, 103].

### v2

@@ 3 Computational Neuroscience Methods for Linguistics
Computational neuroscience provides a methodological bridge between formal linguistic theory and empirical neural data. As illustrated in Fig. 2, it transforms abstract linguistic structures into computational representations that can be quantitatively compared with neural activity through a set of interrelated tools and analytical frameworks. This section introduces the core components of this bridge. We first outline the methodology of computational neuroscience and then describe four technical pillars: Measurement Methods (neural activity recording), Word Embeddings (numerical representations of language), Evolving Language Models, and Neural Coding (analytical frameworks that link model representations to brain signals). Finally, we present an integrative case study to illustrate how these components operate jointly and support model–brain alignment research in linguistics and neuroscience.
@@ 3.2 Brain Activity Measurement for Linguistic Tasks
Testing linguistic theories within a computational framework requires reliable measurements of neural activity during language processing. Research on the neural basis of language therefore employs multiple techniques that capture neural responses across distinct spatial, temporal, and representational scales [99, 75, 236, 113, 84]. Because language processing emerges from dynamic interactions across distributed brain regions and multiple scales, no single method is sufficient. Instead, measurement techniques are selected and combined according to the research focus, such as spatial localization, temporal resolution, or representational specificity.
@@
    The main advantage of fMRI is its high spatial resolution, which enables accurate localization of cortical and subcortical regions involved in language processing, including left perisylvian areas (Broca’s area, Wernicke’s area, supramarginal gyrus, angular gyrus) and subcortical structures such as the caudate nucleus and thalamus. It can also distinguish spatial activation patterns associated with different linguistic levels such as phonological, lexical, semantic, and syntactic processing [18, 195, 19, 241, 256]. Its primary limitation is low temporal resolution: the BOLD response is delayed by approximately 4–8 seconds, with signal peaks typically occurring 6–8 seconds after neuronal activation. Consequently, millisecond-scale processes such as rapid semantic integration cannot be directly captured [80, 154]. In addition, fMRI measurements are sensitive to physiological noise from cardiac and respiratory activity, which complicates experiments involving infants or patients with language disorders. A further limitation is that the BOLD signal is an indirect correlate of neural activity, reflecting hemodynamic changes rather than neuronal firing itself. Overall, fMRI offers high spatial resolution but limited temporal resolution [16].
@@
    EEG (Electroencephalography): EEG records weak electric field fluctuations generated by the synchronous firing of large neuronal populations using scalp electrode arrays [140]. Its key advantage is millisecond-level temporal resolution, which enables precise tracking of rapid neural dynamics during language processing, including responses related to semantic and syntactic analysis. However, electrical signals are attenuated and spatially blurred when the pass through brain tissue, skull, and scalp, resulting in poor spatial localization and significant difficulty in identifying precise neural sources [172, 152]. EEG is therefore highly sensitive to synchronized cortical activity but provides limited spatial precision [149]. This positions EEG as a complement to fMRI, which offers high spatial resolution but limited temporal resolution.
@@
    MEG (Magnetoencephalography): MEG measures weak magnetic fields generated by neuronal electrical activity using highly sensitive sensors – traditionally Superconducting Quantum Interference Devices (SQUIDs) [44, 90] and, more recently, Optically Pumped Magnetometers (OPMs). Synchronous neuronal currents produce magnetic fields that can be detected outside the head, providing an indirect measure of neural activity. Like EEG, MEG offers millisecond temporal resolution and can track rapid processing stages from phonetic perception to syntactic and semantic integration.
    Compared with EEG, MEG generally provides better spatial localization because magnetic fields are minimally distorted by skull and scalp tissues, reducing uncertainties associated with EEG inverse problem [92]. This combination of high temporal resolution and improved (though still limited) spatial accuracy makes MEG particularly suitable for studies requiring precise spatiotemporal characterization of language processes, such as phonetic-to-lexical conversion and syntactic processing [208, 210]. However, MEG is primarily sensitive to tangential cortical sources (e.g., in sulci such as the superior temporal and intraparietal sulci) but is largely insensitive to radial sources (e.g., gyral crowns of the inferior frontal gyrus, angular gyrus, and motor cortex) and to subcortical structures. Source localization also remains an ill-posed inverse problem, and spatial resolution, though better than EEG, remains limited to the order of millimeters to centimeters.
@@
    ERP (Event-Related Potential): ERP is not an independent measurement technique but an analysis method applied to EEG recordings [149]. EEG signals are time-locked to stimulus onset and averaged across repeated presentations, extracting stimulus-related neural responses while reducing ongoing background activity and noise [128]. ERP therefore inherits EEG’s millisecond temporal resolution and provides precise temporal information about the time course of language processing, from early sensory analysis to lexical, semantic, and syntactic integration.
    Many ERP components have been identified in language research, among which N400 and P600 are the most widely studied. The N400 is a negative deflection peaking approximately 400 ms after stimulus onset and is strongly associated with semantic processing; its amplitude increases when words are semantically incongruent or unpredictable in context. The P600 is a positive deflection peaking around 600 ms and is primarily linked to syntactic processing; its amplitude increases in response to syntactic violations or elevated structural complexity [229, 178, 127].
@@
    fNIRS (functional Near-Infrared Spectroscopy): fNIRS is a non-invasive optical imaging technique that measures brain activity through changes in the absorption of near-infrared light by oxygenated and deoxygenated hemoglobin [197, 235]. Near-infrared light emitted from scalp-mounted optodes penetrates superficial cortical tissue and is differentially absorbed by oxyhemoglobin and deoxyhemoglobin. Detectors capture diffusely reflected light, and concentration changes are estimated using the modified Beer–Lambert law, allowing indirect inference of neural activity. Like fMRI, fNIRS relies on hemodynamic responses associated with underlying neural activation.
    fNIRS offers several practical advantages. The equipment is portable and does not require shielded environments, enabling experiments in more natural communication settings. It is also generally less sensitive to motion artifacts than fMRI and, in some contexts, more robust than EEG, making it particularly suitable for studies involving infants, children, or patients with language disorders, as well as for tasks involving overt speech or facial movements [242]. Its temporal resolution lies between fMRI and EEG/MEG, allowing tracking of sub-second hemodynamic changes during tasks such as lexical access and sentence comprehension. Spatial resolution is typically on the order of centimeters, sufficient for localizing broad cortical language regions, though not for resolving fine-grained functional organization [71, 211].
    Limitations include lower spatial resolution than fMRI and limited penetration depth, which precludes reliable measurement of subcortical structures. fNIRS signals are also influenced by individual physiological factors such as scalp thickness, hair properties, and superficial (extracerebral) blood flow, requiring careful experimental design to ensure comparability across participants [137, 191]. Although whole-head coverage is now technically feasible with modular high-density arrays, it remains less common than localized cortical applications due to increased demans on hardware, setup time, and data processing, fNIRS is therefore more suitable for targeted investigations of cortical regions rather than comprehensive brain mapping [211].
@@
    ECoG (Electrocorticography): ECoG is an invasive neurophysiological technique that records electrical potentials directly from the exposed cortical surface using subdural electrode grids or strips placed during surgery [100]. Unlike scalp EEG, ECoG avoids signal attenuation and distortion by the skull and scalp, yielding high signal amplitude, excellent spatial resolution, and sub-millisecond temporal precision [112]. These characteristics make it uniquely capable of resolving fine-grained, region-specific cortical dynamics underlying language processing, including articulatory planning, phonemic encoding, lexical access, and syntactic binding. ECoG is widely used in presurgical mapping for patients with epilepsy or brain tumors, allowing direct localization of essential language areas such as Broca’s region, superior temporal gyrus, and inferior parietal cortex with high clinical and scientific reliability [53]. High-frequency broadband power modulations in ECoG are particularly sensitive to local neuronal population activity and correlate strongly with task-specific language computations. Key limitations include its invasive nature, restricted clinical eligibility, limited spatial coverage determined by electrode placement, and inherent ethical constraints that preclude use in healthy participants.
@@
    SEEG (Stereo‐Electroencephalography): SEEG is an invasive, stereotactic technique that uses multiple thin depth electrodes implanted through small burr holes to record electrical activity from deep brain structures and cortical regions, including those buried in sulci, while traversing white matter tracts. This enables precise three-dimensional mapping of the epileptogenic zone [263]. Unlike ECoG, which samples from the cortical surface, SEEG provides three-dimensional access to both cortical and subcortical targets critical for language control, semantic memory, and pragmatic integration [6]. It combines sub-millisecond temporal resolution with precise three‐dimensional localization, enabling characterization of spatiotemporal dynamics with language networks during both overt and covert processing [269]. SEEG is extensively used in epilepsy presurgical evaluation to map language-related networks and identify epileptogenic zones while preserving essential functions. Similar to ECoG, SEEG is restricted to clinical populations, involves surgical risk, and offers limited whole-brain sampling density [183]. Nevertheless, its unique ability to record from deep and hidden structures makes SEEG irreplaceable for investigating subcortical‐cortical circuits supporting language comprehension, production, and monitoring.
3.3 Word Embeddings
@@
Numerical representations of language provide a critical bridge between linguistics, computer science, and neuroscience. Word embeddings, also referred to as distributed word representations, have become a foundational tool in Natural Language Processing (NLP) and in computational studies of neural language mechanisms [158]. The central idea is to map discrete words onto a continuous, high-dimensional vector space, where semantic similarity is reflected in geometric relations among vectors. This mapping enables symbolic linguistic units to be processed quantitatively by computational models. Unlike traditional one-hot encoding, which merely indicates word identity and captures no semantic relationships, word embeddings represent words through distributed, low-dimensional numerical features learned from distributional patterns in text. These representations capture rich semantic and syntactic associations and have significantly improved performance across a wide range of language tasks, from lexical similarity judgments to machine translation and brain encoding models [93].
@@
Word embeddings are typically learned using machine learning or deep learning models and are grounded in the distributional hypothesis, which holds that word meaning is derived from patterns of contextual usage [207]. Based on training objectives and model architectures, mainstream approaches fall into two broad categories. The first comprises static embedding models such as Word2Vec and GloVe. Word2Vec learns word vectors by predicting local context using either the Skip-gram or CBOW architecture, while GloVe integrates global word-co-occurrence statistics with local context information. Both methods generate fixed vector representations that capture semantic similarity and analogical relations among words [186, 40]. The second category consists of context-dependent embeddings produced by modern Transformer-based pre-trained language models. These models dynamically produce context-sensitive representations, addressing polysemy (a limitation static embeddings cannot resolve) and substantially enhancing semantic representation quality [120, 1]. All such approaches rely on large-scale unlabeled corpora and learn patterns of word co-occurrence that place semantically related words in close proximity within the vector space, providing effective foundations for a wide range of downstream language tasks.
@@
Word embeddings play a central role for three main reasons. First, they connect computational linguistics with neurolinguistics [132]. The distributional principle underlying embeddings is broadly consistent with the brain’s distributed representation of lexical semantics, and similarities in embedding space often correspond to similarities in neural activity patterns measured with techniques such as fMRI and MEG [162]. Understanding these representations therefore facilitates direct comparison between computational and neural semantic structures. Second, embeddings provide the numerical input foundation for modern language models. Natural language is symbolic and discrete, and must be transformed into continuous numerical representations before it can be processed by neural networks. From early static embeddings to the context-dependent representations of LLMs, embedding quality strongly influences model performance [58]. Since later sections examine deep learning models for neural simulation, reviewing embedding principles clarifies how linguistic information enters computational models. Third, research on extending embeddings to complex expressions informs studies of neural representations of higher-level linguistic structure [8]. This work addresses not only lexical processing but also sentence- and discourse-level representations. The strengths and weaknesses of existing extension methods, particularly regarding syntactic preservation and semantic robustness, mirror challenges in modeling hierarchical neural language processing. Reviewing these methods therefore provides conceptual guidance for investigating how the brain encodes complex linguistic structures.
In summary, word embeddings and their extensions provide the core framework for numerical language representation. They have driven rapid progress in NLP while also offering essential tools and theoretical insights for interdisciplinary research on the neural mechanisms of language.
@@ 3.4 The Evolving Language Models
Language models are central tools at the intersection of computational neuroscience and linguistics. By learning statistical regularities from large text corpora, they extract latent semantic and syntactic information and generate outputs that approximate human language conventions. Their core computational objective is to predict lexical sequences based on preceding or surrounding context.
Research on language models is important on both theoretical and practical grounds. Theoretically, their development parallels efforts to understand the neurocognitive mechanisms underlying human language processing. Evaluating how well these models capture contextual dependencies and hierarchical structure provides quantitative benchmarks for studying the computational operations of the brain’s language network. Practically, language models form the backbone of many NLP applications, including machine translation, text generation, semantic analysis. Improving their alignment with human cognition is therefore widely regarded as an important step toward more general and robust artificial intelligence systems.
@@
    PCFG: Probabilistic Context-Free Grammars explicitly model hierarchical syntactic structure and can, in principle, capture certain long-distance dependencies. However, both PCFG and n-gram approaches suffer from limited generalization: they cannot exploit semantic similarity across words or structural similarity across contexts. As a result, they have been largely replaced by neural network models, which learn distributed representations and achieve substantially better generalization performance [110].
@@
    RNNs: Recurrent Neural Networks represent the first generation of neural language models. They encode contextual information by mapping words into embedding spaces and updating hidden states sequentially as each new input is processed [67]. Although RNNs do not explicitly encode hierarchical structure, their recurrent connections allow them, in principle, to capture contextual dependencies, including long-distance relations. In practice, however, RNNs suffer from vanishing gradients, which limits their ability to learn very long-range dependencies. Long Short-Term Memory (LSTM) networks are a widely used variant of RNNs designed to mitigate the vanishing gradient problem through gating mechanisms. This architecture substantially improves long-range dependency modeling and became the dominant approach in early neural language modeling systems [102].
@@
    • Autoregressive LLMs: These models generate text sequentially from left to right by predicting each token based on preceding context. The GPT series is the most prominent example, with parameter scales growing from millions to trillions and training corpora spanning diverse web and literary sources [199, 1, 26, 200, 179]. Such models excel at text generation, dialogue, and long-form continuation tasks. Related systems include ERNIE Bot and the GLM family. Their primary strength lies in fluent generation, although strictly unidirectional modeling limits access to full bidirectional context [228, 63, 267, 103].